Use your knowledge of the financial statement equations and their interrelationships to fill in the missing amounts.8 Each of the folowing independent companles is missing numerical data Required: yaur knowledge of the financial statement equations and their interralationships to fill in the missing amounts. Alpha Co. Bravo, Inc. Charfe Co. S 34,4006400 39,400 61,400 20,700 107,000 (27,200) 132.800 31,400 10.700 61.400 40,700
Answer #1

Company

Total Revenue

Total Expenses

Net income(loss)

Total assets

Total Liabilities

Shareholder's equity

Alpha co

$ 61,400.00

$      40,700.00

$      20,700.00

$ 1,01,800.00

$ 34,400.00

$ 67,400.00

Bravo co

$ 79,800.00

$ 1,07,000.00

$    (27,200.00)

$ 1,32,800.00

$ 93,400.00

$ 39,400.00

Charlie co

$ 31,400.00

$      10,700.00

$      20,700.00

$      61,400.00

$ 40,700.00

$ 20,700.00

Company

Total Revenue

Total Expenses

Net income(loss)

Total assets

Total Liabilities

Shareholder's equity

Alpha co

61400

=61400-20700

20700

=34400+67400

34400

67400

Bravo co

=107000-27200

107000

-27200

132800

=132800-39400

39400

Charlie co

31400

10700

=31400-10700

61400

40700

=61400-40700